OpenStudy (anonymous):

Hey, I'm having a little trouble understanding the last part of the distance formula, if anyone could help me out, that'd be great!

OpenStudy (anonymous):

I get everything except how to get 53, please help me.

OpenStudy (sleepyjess):

53 doesnt have a square root that is a whole number which is why it is left as just \(\sqrt{53}\)

OpenStudy (anonymous):

So if it was different than 53, you just use the square root?

OpenStudy (sleepyjess):

for example if it were \(\sqrt{36}\) it could be simplified to just 6 or \(\sqrt{4}\) would become 2

OpenStudy (sleepyjess):

do you understand what the square root of a number is?

OpenStudy (anonymous):

Yes, what would I do if it became a decimal?

OpenStudy (sleepyjess):

what do you mean? like if the square root was 9.6 for example?

OpenStudy (anonymous):

Yes

OpenStudy (sleepyjess):

then you would just leave it as \(\sqrt{x}\) unless they want you to put the decimal

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (sleepyjess):

\(\Huge\frak\color{magenta}{No~problem!\\If~you~have~any~others~feel~free\\to~tag~me}\) \(\bbox [10pt, blue,border:5pt solid black ]{\color{seagreen}{\huge\cal -sleepyjess}}\)